OpenStudy (anonymous):

how do you solve \sqrt{-10+7p}=p

OpenStudy (anonymous):

-10+7p=p² => then you solve the second order equation p²-7p+10 using delta

OpenStudy (anonymous):

do you have to divide

OpenStudy (anonymous):

p²-7p+10=0 delta=49-40=9 => p1=7-3/2=2 p2=7+3/2=5

OpenStudy (anonymous):

oh ok